# FingerTip 20K: A Benchmark for Proactive and Personalized Mobile LLM Agents

### v1

@@ Abstract
Mobile GUI agents are becoming critical tools for enhancing human-device interaction efficiency, with multimodal large language models (MLLMs) emerging as dominant paradigms in this domain. Current agents, however, are limited to following explicit human instructions, resulting in insufficient capability for proactive intent anticipation. Additionally, these agents fail to leverage the contextual information associated with users during task execution, thereby neglecting potentially vast differences in user preferences. To address these challenges, we introduce the FingerTip benchmark. It contains two new tracks: proactive task suggestions by analyzing environment observation and users’ previous intents, and personalized task execution by catering to users’ action preferences. We collected unique human demonstrations of multi-step Android device interactions across a variety of everyday apps. These demonstrations are not isolated but are continuously acquired from the users’ long-term usage in their real lives, and encompass essential user-related contextual information. Our experiments reveal challenges of the tasks we propose. The model fine-tuned with the data we collected effectively utilized user information and achieved good results, highlighting the potential of our approach in building more user-oriented mobile GUI agents. Our code is open-source at https://anonymous.4open.science/r/FingerTip-57B8 for reproducibility.
@@ 1 Introduction
Recent studies have explored how to utilize multimodal large language models (MLLMs) to construct graphical user interface (GUI) control agents [1, 2, 3, 4, 5], with a significant direction being mobile phone UI control agents, as mobile phone UIs are among the most frequently interacted with by humans in daily life. These agents receive a natural language task instruction, such as "Set an alarm for 7:30 for me", and then perceive the device state by observing the device screen (via screenshots or textual UI trees), and generate actions (click, type, scroll, etc.) to interact with the device environment to fulfill human instructions.
Figure 1: An overview task example in FingerTip.
@@
Despite rapid progress, currently, most existing LLM-based mobile GUI agents are confined to a completely passive paradigm: they only perform tasks upon receiving a clear instruction. This paradigm restricts their ability to proactively offer task suggestions and assistance in the absence of direct human instructions. If users have to formulate detailed instructions for every intent when interacting with mobile GUI agents, it will significantly increase the cognitive burden of mobile phone usage. Moreover, humans sometimes may not clearly express some latent needs. Therefore, mobile GUI agents need to be more proactive to provide users with more comprehensive and seamless services. Furthermore, the existing agents utilize almost exclusively user instructions as textual information when performing tasks, without taking into account any additional user-related information (e.g., time and location, user profile, user historical intents and actions), thus failing to provide personalized services to users.
@@
To comprehensively evaluate the proactive and personalized capabilities of mobile GUI agents, we propose the FingerTip benchmark, which includes two new tracks: (i) proactive task suggestion, where the agent needs to integrate the user’s past intents and the current environmental state to infer the user’s potential current intent; (ii) personalized task execution, where the agent needs to refer to the user’s past action preferences to execute current instructions. Since existing benchmarks do not provide such user-related contextual information, we spent two months collecting new diverse data from 91 users in their daily mobile phone usage, including 21,437 episodes covering 506 apps. We then conducted experiments on the FingerTip benchmark to evaluate the capabilities of generalist models and GUI-control agents built on specifically designed models and found that there is still much room for improvement in their proactive and personalized capabilities. We fine-tuned a small model using the collected data and achieved better results.
@@
  • We propose the FingerTip benchmark, which includes two brand-new tracks, to evaluate the ability of mobile GUI agents to proactively predict user intents and offer suggestions, as well as their ability to personalize task execution in accordance with user preferences.
@@
  • We evaluate the capabilities of generalist models and GUI-control-specific models on the FingerTip benchmark, demonstrating the difficulty of the tracks we propose. The excellent performance of the model fine-tuned with our collected data highlights the potential of our approach in building more proactive and personalized mobile agents.
@@ 2.1 Mobile GUI-control datasets and benchmarks
Table 1 compares FingerTip to existing mobile GUI-control datasets and benchmarks. These datasets typically represent each data instance through two core components: a textual task instruction and its corresponding operational demonstration. The demonstration is encoded as a sequence of interface interactions (e.g., clicking, typing, scrolling) accompanied by relevant screenshots. What differentiates them is mainly whether they are single-step (grounding instructions to UI elements on the screen), and whether they have supplemental View Hierarchy (VH) data for each screenshot. These datasets share some common drawbacks. Firstly, their task instructions are either pre-defined by authors and annotators or generated by LLMs, and it is questionable to what extent they can reflect the real intents of people using mobile phones in their daily lives. Additionally, they collect task demonstrations mainly by having annotators operate simulators or connected phones on computers, which is not the real scenario of people using mobile phones. Finally, each data instance is isolated and does not contain any contextual information related to the user.
Table 1: Comparison of FingerTip to existing mobile GUI-control datasets and benchmarks.
@@
#Collection
method
UI tree? Screen?
Profile
@@
AndroidLab Android 10.5k - - human annotate yes yes no no execution
AMEX Android 8k 110 12.8 human annotate yes yes no no execution
AndroidControl Android 15283 833 5.5 human annotate no yes no no execution
AitW Android 715142 357 6.5 human annotate yes yes no no execution
AndroidLab
benchmark
Android 138 9 8.6
select from
AndroidLab
no yes no no execution
A3 Android 201 20 - unknown yes yes no no execution
SPHINX Android - 100 8.1 human annotate yes yes no no execution
SPA-Bench Android 340 58 - human annotate no yes no no execution
FingerTip Android 21437 506 11.1 human annotate yes yes yes yes
@@
For benchmarks, the success rate is the most commonly used metric, and some studies also consider efficiency and cost. A common approach to assessing the success of a task is to determine whether essential states have been reached [6, 7, 8]. Some studies also compare agents’ actions to golden actions [9]. However, these golden actions do not take into account potentially vast differences in user preferences, that is, the action sequences of different users to complete similar tasks may be very different. In addition, current benchmarks have similar task forms, that is, given an existing instruction, how to perform actions to complete it. To the best of our knowledge, there is no mobile GUI agent benchmark that discusses how to proactively suggest tasks based on user-related information when instructions are unknown.
@@ 2.2 Mobile GUI-control agents
Mobile GUI agents are designed to understand the UI and automate tasks on mobile apps in a manner similar to that of humans. Current agents leverage the extensive world knowledge and powerful embodied capabilities of multimodal large language models (MLLMs) for complex task planning and reasoning in multi-step GUI-control tasks. One notable approach is to directly guide generalist models like GPT-4v to perform tasks through extensive prompt engineering [3, 10, 11, 1, 4, 12]. However, these methods require meticulously designed prompts to achieve the best results. Another research direction focuses on fine-tuning smaller models [13, 14, 15, 16, 17] on GUI-specific datasets to endow them with GUI grounding capabilities and the ability to break down high-level instructions, thereby enhancing their operational efficiency. Despite these advancements, most current agents are still confined to passively following explicit instructions and are unable to proactively predict user needs. Moreover, they do not take into account any user preferences when performing tasks. Some studies focus on proactively clarifying users’ ambiguous instructions [18, 19, 20]; however, these studies still require users to provide initial instructions. Proactive Agent [21] predicts potential tasks by monitoring user activities and environmental states, but the input is text-only, and the task scenarios are mainly limited to computer or web environments rather than mobile ones.
@@ 3.1 Proactive task suggestion
In the FingerTip benchmark we propose, unlike the evaluation tasks of traditional mobile GUI agent benchmarks that rely entirely on explicit instructions, we introduce a new task where the agent proactively predicts the user’s current intent and proposes tasks suggestion that the user might want to perform, as shown in Figure 2. The agent’s task is to generate an intent prediction $I$ based on the user profile $U$, the current time $T$, the current scenario $S$, the user’s historical intents $I_{\text{history}}$, and the partial screenshots $O$ observed at present. This can be formalized as:
@@
where $f$ represents the agent. $I$ is a sentence that unambiguously predicts the intent of the user, including all necessary information. It should clearly state the name of the app that the user wants to use, and the final effect that the user wants to achieve. $U$ includes common user attributes such as age, sex, occupation, etc. $T$ represents the current timestamp, accurate to the second. $S$ represents the current scenario. To avoid revealing real location information, we use some common location categories (such as residence, school, office, etc.) to represent it. $I_{\text{history}}$ contains the user’s historical intents in the recent period, up to 20 items, which may include the potential patterns and preferences of the user’s mobile phone usage. $O$ includes the first few screenshots of the user’s current actions (e.g., opening the home page of a certain app), which should be able to reveal the initial part of the user’s true intent (e.g., Use a certain app to…). It should be noted that we provide at most three such screenshots and at least zero. That is to say, the agent may need to make predictions in the complete absence of $O$. It is obvious that the fewer the number of screenshots, the more difficult it is to predict the user’s true intent. We hope that the agent can utilize the above-mentioned user-related contextual information to infer the user’s potential intents, and thereby proactively offer helpful task suggestions.
@@ 3.2 Personalized task execution
In addition to proactive task suggestion, we also aim to evaluate the agent’s ability to carry out tasks under the condition of explicit instructions, that is, when the user’s intent is known. The setting of this part is similar to the existing benchmarks. The difference lies in that we additionally assess the agent’s ability to execute tasks in a personalized manner specifically catering to the action preferences of different users. Given user profile $U$, user instruction $I_{\text{true}}$, user’s historical actions $A_{\text{history}}$, agent’s action sequence $A_{\text{agent}}$, and the current screenshot $O_{t}$ and the corresponding accessibility tree $AT_{t}$, the agent needs to perform the next action $A_{t+1}$, and then observe $O_{t+1}$ and $AT_{t+1}$. This can be formalized as:
@@
where $f$ represents the agent. $I_{\text{true}}$ is equivalent to the user’s true intent that needs to be predicted in proactive task suggestion, and here it serves as the instruction to be executed. $A_{\text{history}}$ is the complete action sequence of the user when performing a similar task in the past, provided to the agent for in-context learning to imitate the user’s action preferences. $A_{\text{agent}}$, on the other hand, is the action sequence $\{A_{1},...,A_{t}\}$ that the agent has already executed in the current task, helping the agent determine the progress of the task. The agent needs to constantly interact with the mobile phone environment until it believes that the user’s instruction $I_{\text{true}}$ has been fulfilled. We hope that the final sequence of agent actions $A_{\text{agent}}$ can reflect the user’s action preferences.
4 The FingerTip benchmark
@@ 4.1 Overview
The motivation for FingerTip data collection is to evaluate the dual tracks we have proposed, namely proactive task suggestion and personalized task execution. To this end, the most distinctive feature of the data should be user-oriented, containing sufficient user-related contextual information and being able to reflect the patterns and preferences of users in terms of intents and actions.
Figure 3: Data collection pipeline.
@@ 4.2 Data collection
The data collection process is shown in Figure 3. We first recruited 91 data collectors (hereinafter referred to as users) using Android phones through crowdsourcing. Each user filled out a questionnaire, which collected their user profiles. Users were required to download an APP developed by us on their own daily used phones and use it to collect data. Specifically, whenever users had a real intent to use their phones in their daily lives, they could open the FingerTip APP, record their intent at that moment in one sentence, and select the location category they were in. Then, users needed to switch to the app involved in the intent they recorded and demonstrate the specific action sequence to complete this intent.
The FingerTip APP will automatically upload the intent they fill in (including time and location) and the demonstration process they provide (including screenshot sequences, corresponding accessibility tree XML file sequences, and UI action sequences) to the cloud server. This is regarded as the user collecting one piece of data. The APP may remind the user to collect data when they wake up the phone screen to prevent them from forgetting. Each user is required to use their phone to collect data for a minimum of one month and a maximum of two months, with a maximum of 12 pieces of data uploaded per day. In this way, users can fully customize the data they upload. Intents and demonstrations originate from the user’s daily usage scenarios, rather than being pre-defined by researchers or generated by LLMs. Before the formal data collection began, each user underwent training to provide high-quality data. As all users were from mainland China, the data we collected mainly concerned third-party Chinese apps and covered a wide range of device types and Android versions.
FingerTip APP is developed based on the accessibility features of the Android system. It can automatically record the type and coordinates, as well as optional text descriptions of each user action. The actions we collect are unified into an action space, as shown in Table 2. Among them, $terminate()$ is uniformly added to the last screenshot of all episodes. Due to the incomplete development of accessibility features in some Chinese apps, it may be impossible to automatically capture complete action information on certain interfaces of these apps, returning incorrect action formats that do not belong to the action space, or the returned actions lack coordinates. For these non-compliant actions, we use Qwen-VL-Max to observe the two screenshots before and after the action, and regenerate the action, then use PaddleOCR to check if the coordinates are correct.
Figure 4: Dataset statistics and distribution.

### v2

@@ Abstract
Mobile GUI agents are becoming critical tools to improve user experience on smart devices, with multimodal large language models (MLLMs) emerging as the dominant paradigms in this domain. Current agents, however, rely on explicit human instructions, overlooking the potential to leverage the contextual information (like location, time, user profile) and historical data for proactive task suggestions. Besides, previous works focus on optimizing the success rate during task execution, but pay less attention to the personalized execution trajectory, thereby neglecting potentially vast differences in user preferences. To address these challenges, we introduce the FingerTip 20K benchmark. We collected 20K unique human demonstrations of multi-step Android device interactions across a variety of everyday apps. These demonstrations are not isolated but are continuously acquired from the users’ long-term usage in their real lives, and encompass essential user-related contextual information. The benchmark contains two new tracks: proactive task suggestions by analyzing environment observation and users’ previous intents, and personalized task execution by catering to users’ action preferences. Our experiments reveal that the tracks we propose pose significant challenges for leveraging user-related information in GUI tasks. We also performed a human study to show that there exists a huge gap between existing agents and humans. The model fine-tuned with the data we collected effectively utilized user information and achieved good results, highlighting the potential of our approach in building more user-oriented mobile LLM agents. Our code is open-source at https://github.com/tsinghua-fib-lab/FingerTip-20K for reproducibility.
@@ 1 Introduction
Recent studies have explored how to utilize multimodal large language models (MLLMs) to build graphical user interface (GUI) control agents (Koh et al., 2024; Zheng et al., 2024; Yan et al., 2023; Kim et al., 2023; Deng et al., 2023), with a significant direction being mobile phone GUI control agents. These mobile LLM agents have the potential to tremendously improve user experience with mobile devices, since GUI is a universal interface across various applications. These agents receive a natural language task instruction, such as "Set an alarm for 7:30 for me", and then perceive the device state by observing the device screen (via screenshots or textual UI trees), and generate actions (click, type, scroll, etc.) to interact with the device environment to fulfill human instructions.
Figure 1: An overview task example in FingerTip 20K. The agent proactively offers task suggestions to the user and personalizes the execution of tasks in a way that aligns with the user’s preferences.
@@
Despite rapid progress, currently, most existing mobile LLM agents are confined to a completely passive paradigm: they only perform tasks upon receiving a clear instruction. This paradigm restricts their ability to proactively offer task suggestions and assistance in the absence of direct human instructions. If users have to formulate detailed instructions for every intent when interacting with mobile LLM agents, it will significantly increase the cognitive burden of mobile phone usage. Moreover, humans sometimes may not clearly express some latent needs. Therefore, mobile LLM agents need to be more proactive to provide users with more comprehensive and seamless services. Furthermore, the existing agents utilize almost exclusively user instructions as textual information when performing tasks, without taking into account any additional user-related information (e.g., time and location, user profile, user historical intents and actions), thus failing to provide personalized services to users. We argue that these limitations stem largely from the lack of suitable training data and standardized evaluation benchmarks that incorporate rich user-related information.
@@
To comprehensively evaluate the proactive and personalized capabilities of mobile LLM agents, we propose the FingerTip 20K benchmark, which includes two new tracks: (i) proactive task suggestion, where the agent needs to integrate the user’s past intents and the current environmental state to infer the user’s potential current intent; (ii) personalized task execution, where the agent needs to refer to the user’s past action preferences to execute current instructions. The overall task scenario we envision is shown in Figure 1. Since existing benchmarks do not provide user-related contextual information and historical data, we spent over one month collecting new diverse data from 95 users in their daily mobile phone usage, including 21,437 episodes covering 506 apps. We then conducted experiments on the FingerTip 20K benchmark to evaluate the capabilities of generalist models and GUI-control agents built on specifically designed models and found that there is still much room for improvement in their proactive and personalized capabilities. Current agents still find it hard to reach or surpass the human level. The best-performing model achieved a success rate of 12.8%, while humans reached 30.3%. We fine-tuned a small model using the collected data and achieved better results.
@@
  • We propose the FingerTip 20K benchmark, which includes two brand-new tracks, to evaluate the ability of mobile LLM agents to proactively predict user intents and offer suggestions, as well as their ability to personalize task execution in accordance with user preferences.
@@
  • We evaluated the capabilities of generalist models and GUI-control-specific models on the FingerTip 20K benchmark, demonstrating the difficulty of the tracks we propose. The excellent performance of the model fine-tuned with our collected data highlights the potential of our approach in building more proactive and personalized mobile agents.
@@ 2.1 Mobile GUI-control datasets and benchmarks
Table 1 compares FingerTip 20K to existing mobile GUI-control datasets and benchmarks (Chai et al., 2024; Li et al., 2024; Rawles et al., 2023; 2024; Xu et al., 2024a; Chai et al., 2025; Ran et al., 2025; Chen et al., 2024). These datasets typically represent each data instance through two core components: a textual task instruction and its corresponding operational demonstration. The demonstration is encoded as a sequence of interface interactions (e.g., clicking, typing, scrolling) accompanied by relevant screenshots. What differentiates them is mainly whether they are single-step (grounding instructions to UI elements on the screen), and whether they have supplemental View Hierarchy (VH) data for each screenshot. These datasets share some common drawbacks. Firstly, their task instructions are either pre-defined by authors or generated by LLMs, and it is questionable to what extent they can reflect the real intents of people using mobile phones in their daily lives. Additionally, they collect task demonstrations mainly by having annotators operate simulators on computers, which is not the real scenario of people using mobile phones. Finally, each data instance is isolated, lacking temporal correlation and contextual information related to the user.
Table 1: Comparison of FingerTip 20K to existing mobile GUI-control datasets and benchmarks.
@@
User-defined
tasks?
Contextual
@@
Android Instruct 10.5k - 9.0 execution
AMEX 3046 192 12.8 execution
AndroidControl 15283 833 5.5 execution
AitW 715142 357 6.5 execution
AndroidWorld 116 20 - execution
AndroidLab 138 9 - execution
A3 201 20 - execution
SPHINX - 100 8.1 execution
SPA-Bench 340 58 - execution
[…]
@@
For benchmarks, the success rate is the most commonly used metric, and some studies also consider efficiency and cost. A common approach to assessing the success of a task is to determine whether essential states have been reached (Rawles et al., 2024; Zhang et al., 2024; Lee et al., 2024). Some studies also compare agents’ actions to golden actions (Xing et al., 2024). However, these golden actions do not take into account potentially vast differences in user preferences, that is, the action sequences of different users to complete similar tasks may be very different. In addition, current benchmarks have similar task forms, that is, given an existing instruction, how to perform actions to complete it. To the best of our knowledge, there is no mobile LLM agent benchmark that discusses how to proactively suggest tasks based on user-related information when instructions are unknown.
@@ 2.2 Mobile GUI-control agents
Mobile GUI agents are designed to understand the UI and automate tasks on mobile apps in a manner similar to that of humans. Current agents leverage the extensive world knowledge and powerful embodied capabilities of multimodal large language models (MLLMs) for complex task planning and reasoning in multi-step GUI-control tasks. One notable approach is to directly guide generalist models like GPT-4v to perform tasks through extensive prompt engineering (Yan et al., 2023; Rawles et al., 2023; He et al., 2024; Koh et al., 2024; Kim et al., 2023; Zheng et al., 2023; Zhang et al., 2025; Wen et al., 2024). However, these methods require meticulously designed prompts to achieve the best results. Another research direction focuses on fine-tuning smaller models (Nakano et al., 2022; Qin et al., 2025; Hong et al., 2024; Xu et al., 2024b; Gur et al., 2023) on GUI-specific datasets to endow them with GUI grounding capabilities and the ability to break down high-level instructions, thereby enhancing their operational efficiency. Despite these advancements, most current agents are still confined to passively following explicit instructions and are unable to proactively predict user needs. Moreover, they do not take into account any user preferences when performing tasks. Some studies focus on proactively clarifying users’ ambiguous instructions (Wu et al., 2021; Chen et al., 2020; Qian et al., 2024); however, these studies still require users to provide initial instructions. Proactive Agent (Lu et al., 2024) predicts potential tasks by monitoring user activities and environmental states, but the input is text-only, and the task scenarios are mainly limited to computer or web environments rather than mobile ones.
@@ 3.1 Proactive task suggestion
In the FingerTip 20K benchmark we propose, unlike the evaluation tasks of traditional mobile LLM agent benchmarks that rely entirely on explicit instructions, we introduce a new task where the agent proactively predicts the user’s current intent and proposes tasks suggestion that the user might want to perform, as shown in Figure 2. The agent’s task is to generate an intent prediction $I$ based on the user profile $U$, the current time $T$, the current scenario $S$, the user’s historical intents $I_{\text{history}}$, and the partial screenshots $O$ observed at present. This can be formalized as:
@@
where $f$ represents the agent. $I$ is a sentence that unambiguously predicts the intent of the user. It should clearly state the name of the app that the user wants to use, and the final effect that the user wants to achieve. $U$ includes common user attributes such as age, sex, occupation, etc. $T$ represents the current timestamp, accurate to the second. $S$ represents the current scenario, expressed in common location categories. $I_{\text{history}}$ contains the user’s historical intents in the recent period, up to 20 items, which may include the potential patterns and preferences of the user’s mobile phone usage. $O$ includes the first few screenshots of the user’s current actions (e.g., opening the home page of a certain app). We hope that the agent can utilize the above-mentioned user-related contextual information and historical intents to infer the user’s potential intents, and thereby proactively offer helpful task suggestions.
@@ 3.2 Personalized task execution
In addition to proactive task suggestion, we also aim to evaluate the agent’s ability to carry out tasks under the condition of explicit instructions, that is, when the user’s intent is known. The setting of this part is similar to the existing benchmarks. The difference lies in that we additionally assess the agent’s ability to execute tasks in a personalized manner specifically catering to the action preferences of different users. Given user profile $U$, user intent $I_{\text{true}}$, user’s historical actions $A_{\text{history}}$, agent’s action sequence $A_{\text{agent}}$, and the current screenshot $O_{t}$ and the corresponding accessibility tree $AT_{t}$, the agent needs to perform the next action $A_{t+1}$, and then observe $O_{t+1}$ and $AT_{t+1}$. This can be formalized as:
@@
where $f$ represents the agent. $I_{\text{true}}$ is equivalent to the user’s true intent that needs to be predicted in proactive task suggestion, and here it serves as the instruction to be executed. $A_{\text{history}}$ is the complete action sequence of the user when performing a similar task in the past, provided to the agent for in-context learning to imitate the user’s action preferences. $A_{\text{agent}}$, on the other hand, is the action sequence $\{A_{1},...,A_{t}\}$ that the agent has already executed in the current task, helping the agent determine the progress of the task. The agent needs to constantly interact with the mobile phone environment until it believes that $I_{\text{true}}$ has been fulfilled. We hope that the final sequence of agent actions $A_{\text{agent}}$ can reflect the user’s action preferences.
4 The FingerTip 20K benchmark
@@ 4.1 Overview
The motivation for FingerTip 20K data collection is to evaluate the dual tracks we have proposed, namely proactive task suggestion and personalized task execution. To this end, the most distinctive feature of the data should be user-oriented, containing sufficient user-related contextual information and being able to reflect the patterns and preferences of users in terms of intents and actions.
Figure 3: Data collection pipeline. Users record their intents and demonstrate actions by using the FingerTip APP in their daily mobile phone usage.
@@ 4.2 Data collection
The data collection pipeline is shown in Figure 3. We first recruited 95 data collectors (hereinafter referred to as users) using Android phones through crowdsourcing, covering a wide range of device types and Android versions. Users were required to download an APP developed by us on their own daily used phones and use it to collect data. Specifically, whenever users had a real intent to use their phones in their daily lives, they could open the FingerTip APP, record their intent at that moment in one sentence, and select the location category they were in. Then, users needed to switch to the app involved in the intent they recorded and demonstrate the specific action sequence to complete this intent.
The FingerTip APP will automatically upload the intent they fill in (including time and location) and the demonstration process they provide (including screenshot sequences, corresponding accessibility tree XML file sequences, and UI action sequences) to the cloud server. This is regarded as the user collecting one piece of data. The APP may remind the user to collect data when they wake up the phone screen to prevent them from forgetting. Each user is required to use their phone to collect data for one month, with a maximum of 12 pieces of data uploaded per day. In this way, users can fully customize the data they upload. See Appendix A.3 and A.4 for more details on data collection and data format.
FingerTip APP is developed based on the accessibility features of the Android system. It can automatically record the type and coordinates, as well as optional text descriptions of each user action. The actions we collect are unified into an action space, as shown in Table 2. Among them, finishfinish is uniformly added to the last screenshot of all episodes.
Figure 4: Dataset statistics and distribution. (a) The length distribution of the natural language intents recorded by users. (b) The distribution of the number of screenshots contained in each episode (i.e., the distribution of the number of action steps of users). (c) The distribution of all categories to which the intents belong. (d) The distribution of all apps involved in the data.